(13/4x) + (3/x^2) = (5/2x); Solve for x
Answer by DRichardson(63)

(13/4x) + (3/x^2) = (5/2x); Solve for x
Hello There,
x = -(sqrt(3)*4^(1/3)*%i-4^(1/3))/2

solution 2

x = (sqrt(3)*4^(1/3)*%i+4^(1/3))/2


solution 3

x = -4^(1/3)

Hope this helps!
Delilah Richardson